The sum of the roots of equation 5x² -10x + 3 =0
 
.......

Grade:12

2 Answers

Jitendra Kumar
3 years ago
5x2 -10x +3 = 0
We know that general form of quadratic equation ax2 + bx + c = 0
Now comparing both equation so we get 
a=5, b=-10 and c=3
We also know that
Sum of the root = -b/a
So sum of the root = -(-10)/5 = 2
Hence, Sum of the root = 2
Ram Kushwah
3 years ago
The given equation is:
 
5x²-10x+3=0...........................(1)
 
This is in the form :
ax²+bx+c=0
here we no that sum of roots=-b/a
In the given eqn (1)
a=5 and b=-10
Thus sum of roots =- (-10/5)
=-(-2)=2
Hence sum of roots= 2
Thanks 
Pl aprove